Calculate the wavelength, frequency and wavenumber of a light wave whose period is 2.0 × 10–10 s.

Grade:12

1 Answers

Gaurav
9 years ago
Hello Student
Frequency = 1/Time Period = 5 x 109 s-1
Wavelength =c/v = 6.10-2m
Wave number = 1/wavelength = 16.55 m